If the inverse demand function for toasters is p=70-Q, what is the consumer surplus if price is $25? The consumer surplus is (round your answer to two decimal places)
Answer #1

Answer
Consumer surplus=0.5*(maximum willingness to pay -price)*Q
P=25
25=70-Q
Q=45
CS=0.5*(70-25)*45
=1012.50
The CS is $1012.50